# Light-Weight Fault Tolerant Attention for Large Language Model Training

### v2

@@ 4.3. Correcting Extreme Error Propagations
Next we describe how we adaptively handle three common error patterns in the attention mechanism and discuss how we efficiently parallelize the correction procedures on GPUs. Note that for all cases, we limit EEC-ABFT to handle up to 1D patterns since 2D patterns cannot be efficiently corrected. We defer our discussion on propagation mitigation to Section 5 where we design systematic ABFT schemes for attention.
@@
Deterministic Patterns: An error pattern is deterministic for an operation when only one type of propagation pattern can occur in its output. In this case, we only need to maintain the corresponding checksums for error-correcting. For example, 1C pattern only needs row checksums and 1R pattern only needs column checksums. 0D pattern can be corrected using either checksum. Figure 4 shows error-correcting on two matrices: $A$ and $C$. Assuming $A$ has a deterministic 1R error, we illustrate how to detect such patterns efficiently in parallel using EEC-ABFT. Specifically, we let each thread execute EEC-ABFT on column vector in parallel on GPU. We use $T_{0}$ $T_{3}$ to represent four threads for correcting $A$ as an example. For fault-free execution, all executions are divergence-free on GPUs. For 0D pattern, only the error-handling thread will execute the correction procedure and all other threads will exit early to release the GPU resources. For 1D patterns, all threads will execute the correction procedure divergent-free when all error types are the same. Slight divergence may occur when a mixed-type pattern occurs.
@@
Figure 4. Deterministic (1R in $A$) and Nondeterministic (1C in $C$) Error Pattern Handling Using EEC-ABFT
@@
Nondeterministic Patterns: Nondeterministic patterns raise two challenges: (1) the pattern is unknown in advance; (2) when corrupted elements are used for the following operation, they may not only propagate the error to the output but also corrupt the checksums of the output. For 1C pattern, column checksums will be corrupted. For 1R pattern, on the other hand, the row checksums will be corrupted. Figure 4 illustrates the process of recovering a nondeterministic 1D error with 1C being the true pattern and column checksum being corrupted (due to $C^{c}=A^{c}\times B$, assuming $A$ is error free after correction). To handle such a pattern, we need to maintain both column and row checksums. To recover, we first try to use column checksums to fix any potential errors. We will encounter one the two cases: (1) If non-extreme errors occur: Since the column checksums are calculated using the corrupted value, they will wrongly hold a relationship with the corrupted $C$, so no error is detected (false negative). (2) if extreme errors occur: EEC-ABFT will recognize and report the propagation (case 4 of Section 4.2). In either case, no error correction will be done and we will then proceed to use uncorrupted row checksums to fix the errors. Finally, we will use the corrected columns to recover the corrupted column checksums using re-computation. If 1R pattern occurs, on the other hand, column checksums would be able to correct all errors in the first step.
@@
Mixed-type Patterns: When a mixed-type pattern occurs, we cannot directly conclude the error type based on the value $\delta_{1}$, because of error type transition. For example, near-INF in the 1D pattern can lead to $\delta_{1}$ being INF, which may cause confusion for error detection. EEC-ABFT overcomes such a challenge by counting all possible error types before making conclusions error pattern. For example, if $\delta_{1}=INF$, we count both INF and near-INF. If $\delta_{1}=NAN$, all three kinds of error need to be considered.
@@ 4.4. Systematic ABFT Scheme for Attention
Next, we describe how to use EEC-ABFT to build systematic ABFT protection for attention. As our EEC-ABFT can correct errors up to 1D, we divide the whole execution flow of attention mechanism into three protection sections to minimize the probability of more serious error propagations (e.g., 2D). Specifically, the three sections are: $S_{AS}=\{X\times W^{Q},X\times W^{K},Q\times K^{T}\}$, $S_{CL}=\{X\times W^{V},AP\times V\}$, $S_{O}=\{CL\times W^{O}\}$. In addition, to ensure propagated errors can be effectively corrected and reduce error detection overhead, we carefully design checksum passing mechanism for each section. Specifically, we only need to compute the initial checksums for the inputs of the first operation and pass the updated checksums between operations to form protection for a whole section. This enables us to correct up to 1 error at a time per section.
@@
Attention Score Protection Section: As shown in Figure 5(a), to protect the calculation of the attention score, we encode the input matrix $X$ with column checksums. After the multiplying input with parameter matrices, the output $Q$ and $K$ matrix will have their corresponding column checksums. If any computational error occurs during the multiplication, it will manifest as 0D pattern shown as red/green element in $Q$ and $K$. Although we can use their column checksums to correct those errors, we choose to lower the error handling overhead by delaying detecting and correcting until $Q\times K^{T}$ is done. Pre-existing errors in $Q$ or $K$ will propagate as a 1D pattern (shown as a red/green column in $AS$), which is tolerable using EEC-ABFT. Finally, 0D errors that occur during $Q\times K^{T}$ (shown as a blue element in $AS$) can also be corrected immediately after computing $AS$.
Context Layer Protection Section: Figure 5(b) shows our protection mechanism for the calculation of $CL$. We encode the parameter matrix $W^{V}$ with row checksums. By doing so, $V$ will have row checksums encoded after $X\times W^{V}$. Next, we encode $AP$ with column checksum, so that after multiplying $AP$ with $V$, the resulting $CL$ will have both row and column checksums. If errors occur during $X\times W^{V}$, it will manifest as a 0D error pattern shown as a yellow element in $V$. Similarly, we delay the error detection until $CL$ is computed. Therefore, any 0D error in $V$ will manifest as a 1D error in $CL$, which can be corrected using its checksums. In addition, 0D errors that occur during $AP\times V$ (shown as purple elements in $CL$) can be corrected using the checksums of $CL$.
@@ 4.5. Adaptive ABFT Detection Frequencies
Based on our ABFT scheme design, each segment can tolerate up to one error that occurs in any operation. As an error does not always lead to non-trainable states, we quantify the vulnerability of an operation by defining $\phi_{OP}^{T}$ being the probability of a type $e$ error in $OP$ that leads to a non-trainable state, which can be estimated by profiling (Table 3). Assuming $f_{S}$ represents the frequency at which ABFT is enabled for the entire series $S$ (e.g., $f_{S}$ = 0.5 corresponds to enabling ABFT every other time we execute $S$), we define $H^{e}_{i}=f_{S}+(1-f_{S})\phi_{OP_{i}}^{e}$ This expression represents the probability of encountering a type $e$ error that can be either correctly handled by ABFT or not handled but does not result in a non-trainable state. So, we define a quantified estimation of the probability that ABFT can correct all errors in $S$ fault coverage (FC).
@@ 4.6. Performance Optimization for GPUs
Encoding: Enabling ABFT for batched matrix operations (e.g., $K\times V$ ) creates an unique computation pattern. This pattern is not well-optimized in vendor-optimized libraries such as NVIDIA cuBLAS. To overcome this, we design highly customized and efficient kernels for encoding checksums. In our kernel, we parallelize the encoding process along the Streaming Multiprocessor by the Number of Heads $\times$ Number of Batches to improve GPU occupancy. To encode each block, we load all data into shared memory and decoupled thread-data mapping between loading and computing (Chen et al., 2021, 2019) such that we can achieve fully coalesced global memory accesses while minimizing bank conflict in shared memory.
@@ 5.4. Adaptive ABFT Detection Frequencies
To demonstrate how ATTNChecker can be adaptive to various error rates representing a wider spectrum of systems, we use our adaptive ABFT optimization algorithm to optimize ABFT detection frequencies for the three detection sections: $S_{AS}$, $S_{CL}$, and $S_{O}$. We assume the system error rate varies from 13 to 20 errors per $10^{25}$ flops for all three types of error according to the latest field report (Dubey et al., 2024). Our target fault coverage is 1 failure per $10^{11}$ executions of attention mechanism based on the operation vulnerability of the Bert model. Figure 10 shows the per-step training overhead with optimized ABFT checking frequencies. When the error rate is small, we can see that most of the ABFT overhead can be reduced (even to 0 when ABFT is not needed). As the error rate increases, ABFT overhead steadily increase with the detection frequencies yet remains minimal (less than 3.6%).
@@ 5.5. Recovery Overhead
In addition to detection, error correction brings extra overhead. Our evaluation show that the correcting 1D error cause by error propagation from $Q$, $K$, and $V$ incurring $0.7\%$ overhead on average. 0D errors can be corrected with 0.3% overhead on average. Correcting errors in $O$ brings 3.9% overhead on average since unlike matrices such as $AS$ where error correction is done within a small block, $O$ is a larger matrix merged from smaller blocks, which requires potentially higher recovery cost.
@@ 6. Related Works
Algorithm-Based Fault Tolerance (ABFT) techniques, on the other hand, are tailored to specific algorithms, leveraging their inherent properties to detect and correct various types of errors, including soft errors, without requiring extensive modifications to the underlying hardware or software. ABFT has been applied to protect and handle errors in matrix decomposition (Huang and Abraham, 1984; Wu and Chen, 2014; Wu et al., 2016), and matrix multiplications (Chen and Dongarra, 2008; Chen et al., 2016a, 2018, b). As ABFT can effectively protect matrix multiplications, several works have extended it to protect ML models such as CNN (Zhao et al., 2020) and Deep learning recommendation systems (Li et al., 2022). ATTNChecker also leverages ABFT but focuses on more computationally-intensive and model-quality-critical attention computation in LLMs.
Resilience of ML models under transient faults. Understanding and evaluating the vulnerability of ML models to hardware faults is becoming increasingly important as ML models scale (Dixit et al., 2022). Existing works investigate various tools and metrics to identify, mitigate, and quantify the impact of faults in ML systems. Existing work investigate various tools and metrics to identifying, mitigating, and quantifying the impact of faults in ML systems. Ares (Reagen et al., 2018) and PyTorchFI (Mahmoud et al., 2020) are fault injection tools build on PyTorch to help quantifying model accuracy drop with faults in ML models. Li et al. (Li et al., 2017) and Agarwal et al. (Agarwal et al., 2023) investigate the Silent Data Corruption (SDC) rate for CNNs and LLMs, respectively, using fault injection. While these works provide valuable insights into the resilience of ML models under transient faults, they primarily focus on fault injection and quantifying the impact of faults. In contrast, ATTNChecker offers a lightweight, real-time solution for not only detecting but also correcting errors during LLM training.
@@ 7. Conclusion
The growing complexity and size of LLMs demand effective fault tolerance to maintain efficient and reliable training. Traditional fault tolerance mechanism such as checkpointing are resource-intensive and time-consuming, hindering the efficiency of LLM training. In response, we introduce ATTNChecker, an algorithm-based fault tolerance technique which offers a lightweight, real-time solution for detecting and correcting errors during LLM training. ATTNChecker significantly reduces recovery overheads by up to 49$\times$ compared to traditional methods, while only adding negligible overhead to the training process. By integrating ATTNChecker into the PyTorch library and testing on various LLMs, we demonstrate its efficacy in enhancing the resilience and efficiency of LLM training. This advancement provides a robust framework for future fault tolerance in large-scale machine learning models.

### v3

@@ 4.3. Correcting Extreme Error Propagations
Next, we describe how we adaptively handle three common error patterns in the attention mechanism and discuss how we efficiently parallelize the correction procedures on GPUs. Note that we limit EEC-ABFT to handle up to 1D patterns for all cases since 2D patterns cannot be efficiently corrected. We defer our discussion on propagation mitigation to Section 5, where we design systematic ABFT schemes for attention.
@@
Deterministic Patterns: An error pattern is deterministic for an operation when only one type of propagation pattern can occur in its output. In this case, we only need to maintain the corresponding checksums for error correcting. For example, the 1C pattern only needs row checksums, and the 1R pattern only needs column checksums. The 0D pattern can be corrected using either checksum. Figure 4 shows error-correcting on two matrices: $A$ and $C$. Assuming $A$ has a deterministic 1R error, we illustrate how to detect such patterns efficiently in parallel using EEC-ABFT. Specifically, we let each thread execute EEC-ABFT on column vector in parallel on GPU. We use $T_{0}$ $T_{3}$ to represent four threads for correcting $A$ as an example. For fault-free execution, all executions are divergence-free on GPUs. For the 0D pattern, only the error-handling thread will execute the correction procedure, and all other threads will exit early to release the GPU resources. For 1D patterns, all threads will execute the correction procedure divergent-free when all error types are the same. Slight divergence may occur when a mixed-type pattern occurs.
@@
Figure 4. Deterministic (1R in $A$) and Nondeterministic (1C in $C$) Error Pattern Handling Using EEC-ABFT ($T_{0}$-$T_{3}$ represents four threads)
@@
Nondeterministic Patterns: Nondeterministic patterns raise two challenges: (1) the pattern is unknown in advance; (2) when corrupted elements are used for the following operation, they may not only propagate the error to the output but also corrupt the checksums of the output. For the 1C pattern, column checksums will be corrupted. For the 1R pattern, on the other hand, the row checksums will be corrupted. Figure 4 illustrates the process of recovering a nondeterministic 1D error with 1C being the true pattern and column checksum being corrupted (due to $C^{c}=A^{c}\times B$, assuming $A$ is error-free after correction). We need to maintain both column and row checksums to handle such a pattern. To recover, we first try using column checksums to fix potential errors. We will encounter one of two cases: (1) If non-extreme errors occur: Since the column checksums are calculated using the corrupted value, they will wrongly hold a relationship with the corrupted $C$, so no error is detected (false negative). (2) if extreme errors occur: EEC-ABFT will recognize and report the propagation (case 4 of Section 4.2). In either case, no error correction will be done and we will then proceed to use uncorrupted row checksums to fix the errors. Finally, we will use the corrected columns to recover the corrupted column checksums using re-computation. If a 1R pattern occurs, on the other hand, column checksums would be able to correct all errors in the first step.
@@
Mixed-type Patterns: When a mixed-type pattern occurs, we cannot directly conclude the error type based on the value $\delta_{1}$, because of error type transition. For example, near-INF in the 1D pattern can lead to $\delta_{1}$ being INF, which may cause confusion for error detection. EEC-ABFT overcomes such a challenge by counting all possible error types before making conclusions about error patterns. For example, if $\delta_{1}=INF$, we count both INF and near-INF. If $\delta_{1}=NAN$, all three kinds of error need to be considered.
@@ 4.4. Systematic ABFT Scheme for Attention
Next, we describe how to use EEC-ABFT to build systematic ABFT protection for attention. As our EEC-ABFT can correct errors up to 1D, we divide the whole execution flow of attention mechanism into three protection sections to minimize the probability of more serious error propagations (e.g., 2D). Specifically, the three sections are: $S_{AS}=\{X\times W^{Q},X\times W^{K},Q\times K^{T}\}$, $S_{CL}=\{X\times W^{V},AP\times V\}$, $S_{O}=\{CL\times W^{O}\}$. In addition, to ensure propagated errors can be effectively corrected and reduce error detection overhead, we carefully design a checksum passing mechanism for each section. Specifically, we only need to compute the initial checksums for the inputs of the first operation and pass the updated checksums between operations to form protection for a whole section. This enables us to correct up to 1 error at a time per section.
@@
Attention Score Protection Section: As shown in Figure 5(a), to protect the calculation of the attention score, we encode the input matrix $X$ with column checksums. After the multiplying input with parameter matrices, the output $Q$ and $K$ matrix will have their corresponding column checksums. If any computational error occurs during the multiplication, it will manifest as a 0D pattern shown as a red/green element in $Q$ and $K$. Although we can use their column checksums to correct those errors, we reduce the error handling overhead by delaying detecting and correcting until $Q\times K^{T}$ is done. Pre-existing errors in $Q$ or $K$ will propagate as a 1D pattern (shown as a red/green column in $AS$), which is tolerable using EEC-ABFT. Finally, 0D errors that occur during $Q\times K^{T}$ (shown as a blue element in $AS$) can also be corrected immediately after computing $AS$.
Context Layer Protection Section: Figure 5(b) shows our protection mechanism for the calculation of $CL$. We encode the parameter matrix $W^{V}$ with row checksums. By doing so, $V$ will have row checksums encoded after $X\times W^{V}$. Next, we encode $AP$ with column checksum so that after multiplying $AP$ with $V$, the resulting $CL$ will have both row and column checksums. If errors occur during $X\times W^{V}$, it will manifest as a 0D error pattern shown as a yellow element in $V$. Similarly, we delay the error detection until $CL$ is computed. Therefore, any 0D error in $V$ will manifest as a 1D error in $CL$, which can be corrected using its checksums. In addition, 0D errors that occur during $AP\times V$ (shown as purple elements in $CL$) can be corrected using the checksums of $CL$.
@@ 4.5. Adaptive ABFT Detection Frequencies
Based on our ABFT scheme design, each segment can tolerate up to one error that occurs in any operation. As an error does not always lead to non-trainable states, we quantify the vulnerability of an operation by defining $\phi_{OP}^{T}$ being the probability of a type $e$ error in $OP$ that leads to a non-trainable state, which can be estimated by profiling (Table 4). Assuming $f_{S}$ represents the frequency at which ABFT is enabled for the entire series $S$ (e.g., $f_{S}$ = 0.5 corresponds to enabling ABFT every other time we execute $S$), we define $H^{e}_{i}=f_{S}+(1-f_{S})\phi_{OP_{i}}^{e}$. This expression represents the probability of encountering a type $e$ error that can be either correctly handled by ABFT or not handled but does not result in a non-trainable state. So, we define a quantified estimation of the probability that ABFT can correct all errors in $S$ fault coverage (FC).
@@ 4.6. Performance Optimization for GPUs
Encoding: Enabling ABFT for batched matrix operations (e.g., $K\times V$ ) creates an unique computation pattern. This pattern is not well-optimized in vendor-optimized libraries such as NVIDIA cuBLAS. We design highly customized and efficient kernels for encoding checksums to overcome this. In our kernel, we parallelize the encoding process along the Streaming Multiprocessor by the Number of Heads $\times$ Number of Batches to improve GPU occupancy. To encode each block, we load all data into shared memory and decoupled thread-data mapping between loading and computing (Chen et al., 2021, 2019) such that we can achieve fully coalesced global memory accesses while minimizing bank conflict in shared memory.
@@ 5.4. Adaptive ABFT Detection Frequencies
To demonstrate how ATTNChecker can be adaptive to various error rates representing a wider spectrum of systems, we use our adaptive ABFT optimization algorithm to optimize ABFT detection frequencies for the three detection sections: $S_{AS}$, $S_{CL}$, and $S_{O}$. We assume the system error rate varies from 13 to 20 errors per $10^{25}$ flops for all three types of error according to the latest field report (Dubey et al., 2024). Our target fault coverage is 1 failure per $10^{11}$ executions of attention mechanism based on the operation vulnerability of the Bert model. Figure 10 shows the per-step training overhead with optimized ABFT checking frequencies. When the error rate is small, we can see that most of the ABFT overhead can be reduced (even to 0 when ABFT is not needed). As the error rate increases, ABFT overhead steadily increases with the detection frequencies yet remains minimal (less than 3.6%).
@@ 5.5. Recovery Overhead
In addition to detection, error correction brings extra overhead. Our evaluation shows that the correcting 1D error caused by error propagation from $Q$, $K$, and $V$ incurs $0.7\%$ overhead on average. 0D errors can be corrected with 0.3% overhead on average. Correcting errors in $O$ brings 3.9% overhead on average since unlike matrices such as $AS$ where error correction is done within a small block, $O$ is a larger matrix merged from smaller blocks, which requires potentially higher recovery cost.
@@ 6. Related Works
Algorithm-Based Fault Tolerance (ABFT) techniques, on the other hand, are tailored to specific algorithms, leveraging their inherent properties to detect and correct various types of errors, including soft errors, without requiring extensive modifications to the underlying hardware or software. ABFT has been applied to protect and handle errors in matrix decomposition (Huang and Abraham, 1984; Wu and Chen, 2014; Wu et al., 2016), and matrix multiplications (Chen and Dongarra, 2008; Chen et al., 2016a, 2018, b). As ABFT can effectively protect matrix multiplications, several works have extended it to protect ML models such as CNN (Zhao et al., 2020) and Deep learning recommendation systems (Li et al., 2022). ATTNChecker also leverages ABFT but focuses on more computationally intensive and model-quality-critical attention computation in LLMs.
Resilience of ML models under transient faults. Understanding and evaluating the vulnerability of ML models to hardware faults is becoming increasingly important as ML models scale (Dixit et al., 2022). Existing work investigates various tools and metrics to identify, mitigate, and quantify the impact of faults in ML systems. Ares (Reagen et al., 2018) and PyTorchFI (Mahmoud et al., 2020) are fault injection tools built on PyTorch to help quantify model accuracy drop with faults in ML models. Li et al. (Li et al., 2017) and Agarwal et al. (Agarwal et al., 2023) investigate the Silent Data Corruption (SDC) rate for CNNs and LLMs, respectively, using fault injection. While these works provide valuable insights into the resilience of ML models under transient faults, they primarily focus on fault injection and quantifying the impact of faults. In contrast, ATTNChecker offers a lightweight, real-time solution for not only detecting but also correcting errors during LLM training.
@@ 7. Conclusion
The growing complexity and size of LLMs demand effective fault tolerance to maintain efficient and reliable training. Traditional fault tolerance mechanisms such as checkpointing are resource-intensive and time-consuming, hindering the efficiency of LLM training. In response, we introduce ATTNChecker, an algorithm-based fault tolerance technique that offers a lightweight, real-time solution for detecting and correcting errors during LLM training. ATTNChecker significantly reduces recovery overheads by up to 49$\times$ compared to traditional methods, while only adding negligible overhead to the training process. By integrating ATTNChecker into the PyTorch library and testing on various LLMs, we demonstrate its efficacy in enhancing the resilience and efficiency of LLM training. This advancement provides a robust framework for future fault tolerance in large-scale machine learning models.
Acknowledgment
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, ComPort: Rigorous Testing Methods to Safeguard Software Porting, under Award Number 78284. The platforms used for evaluation in this work are supported by by the U.S. DOE Office of Science, Office of Advanced Scientific Computing Research, under award 66150: ”CENATE - Center for Advanced Architecture Evaluation”. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under Contract DE-AC05-76RL01830.
Appendix A Artifact Appendix
This artifact contains instructions for how to replicate the experiments. It is available at: zenodo.org/records/14503617.
A.1. Evaluation environment
  • Anaconda virtual environment with python 3.8.10
  • CUDA Toolkit 12.6
  • gcc 11.4.0
  • Nividia A100 - 80GB
A.2. Build and Config
Choose one of the two build options
A.2.1. Option 1: Docker Image
Pull and run a pre-built docker image from Docker Hub.
\$ docker pull lyh911/attnchk-pytorch:2.0
\$ docker run --ipc=host --shm-size=512m --gpus all -it --rm lyh911/attnchk-pytorch:2.0
A.2.2. Option 2: Build from source
1. Create an Anaconda environment.
\$ conda create --name attnchk python==3.8.10
\$ conda activate attnchk
2. Download the source code.
\$ git clone https://github.com/liangyh911/ATTNChecker.git
\$ cd ATTNChecker
3. Install the required Python packages for ATTNChecker.
\$ pip install -r requirements.txt
4. Move the modeling scripts to Huggingface Transformers package.
\$ cp ./HF-moding/modeling_bert.py <Path_To_Anaconda>/envs/attnchk/lib/python3.8/site-packages/transformers/models/bert/modeling_bert.py
\$ cp ./HF-moding/modeling_gpt2.py <Path_To_Anaconda>/envs/attnchk/lib/python3.8/site-packages/transformers/models/gpt2/modeling_gpt2.py
\$ cp ./HF-moding/modeling_gpt_neo.py <Path_To_Anaconda>/envs/attnchk/lib/python3.8/site-packages/transformers/models/gpt_neo/modeling_gpt_neo.py
\$ cp ./HF-moding/modeling_roberta.py <Path_To_Anaconda>/envs/attnchk/lib/python3.8/site-packages/transformers/models/roberta/modeling_roberta.py
5. Build PyTorch from the source.
\$ cd ./ATTNChecker/pytorch
\$ conda install cmake ninja
\$ pip install mkl-static mkl-include
\$ pip install -r requirements.txt
\$ git submodule sync
\$ git submodule update --init --recursive
\$ export CMAKE_PREFIX_PATH=\${CONDA_PREFIX:-"\$(dirname\$(whichconda))/../"}
\$ python setup.py develop
If Python builds successfully, you will see in the last line of the output:
Finished processing dependencies for torch==2.3.0a0+git83e4d29
Test if CUDA is available by using the following codes. $True$ will be returned if cuda is available.
\$ import torch
\$ torch.cuda.is_available()
A.3. Measuring ATTNChecker running overhead
Use .$/records/cleanRecords.py$ to clean the existing records before running new scripts.
\$ cd ./ATTNChecker
\$ python ./records/cleanRecords.py
To measure the overhead of ATTNChecker, use the scripts in .$/ABFT\_running\_time$ folder. The default settings are
  • batch size: 8
  • number of training: 20
Before running the scripts, make sure to use one GPU if you have multiple GPUs on your device. Using export command.
\$ export CUDA_VISIBLE_DEVICES=0
Clean the previous records before running each test.
# bert
\$ python ./records/cleanRecords.py
\$ python ./ABFT_running_time/bertTest.py
# gpt2
\$ python ./records/cleanRecords.py
\$ python ./ABFT_running_time/gpt2.py
# gpt neo
\$ python ./records/cleanRecords.py
\$ python ./ABFT_running_time/gpt-neo.py
# roberta
\$ python ./records/cleanRecords.py
\$ python ./ABFT_running_time/roberta.py
Here is an example output of the BERT model. For each test, the output results may vary. (The loss here is the 1-step training loss.)
Attention Mechanism Overhead: 0.14775651309451615
Training Overhead: 0.0552227860653921
ATTNChecker Loss: 0.5106
no ATTNChecker Loss: 0.5106
A.4. Measuring checkpoint save and load overhead
Before measuring the save and load overhead of the checkpoint, please ensure you have tested ATTNChecker Running Overhead.
Use the scripts in .$/Checkpoint\_time$ folder to test the save and load time of checkpointing a model. Here is an example.
# bert
\$ python ./records/cleanRecords.py
\$ python ./Checkpoint_time/bert.py
# gpt2
\$ python ./records/cleanRecords.py
\$ python ./Checkpoint_time/gpt2.py
# gpt neo
\$ python ./records/cleanRecords.py
\$ python ./Checkpoint_time/gpt-neo.py
# roberta
\$ python ./records/cleanRecords.py
\$ python ./Checkpoint_time/roberta.py
Here is an example output of the BERT model. For each test, the output results may vary. The overhead is calculated in the same way as ATTNChecker running overhead.
Overhead of Checkpointing: 8.522816600251735
A.5. Measuring training loss of ATTNChecker and baseline during 3 epochs
# bert
\$ python ./records/cleanRecords.py
\$ python ./ABFT_epoch_loss/bert.py
# gpt2
\$ python ./records/cleanRecords.py
\$ python ./ABFT_epoch_loss/gpt2.py
# gpt neo
\$ python ./records/cleanRecords.py
\$ python ./ABFT_epoch_loss/gpt-neo.py
# roberta
\$ python ./records/cleanRecords.py
\$ python ./ABFT_epoch_loss/roberta.py
Here is an example output of the BERT model. For each test, the output results may vary. The baseline is training without ATTNChecker.
Loss of ATTNChecker:
1st epoch: 0.5349 , 2nd epoch: 0.3071 , 3rd epoch: 0.1285
Loss of Baseline:
1st epoch: 0.5635 , 2nd epoch: 0.3362 , 3rd epoch: 0.1312